OpenStudy (anonymous):

Order 34 x 10^2, 1.2 x 10^7, 8.11 x 10^-3, and 435 from least to greatest.

OpenStudy (anonymous):

This is my last question of the night.. someone please help me!

OpenStudy (anonymous):

\[34 \times 10^2 = 3400\] \[1.2 \times 10^7 = 12000000\] \[8.11 \times 10^{-3} = 0.00811\] \[435 = 435\] Now, can you solve further or not??

OpenStudy (anonymous):

Yeah I got it thanks!

OpenStudy (anonymous):

Welcome dear...